Martinez Inc. presented the following data. Net income $2,550,000 Preferred stock: 51,000 shares outstanding, $100 par, 8% cumulative, not convertible 5,100,000 Common stock: Shares outstanding 1/1 816,000 Issued for cash, 5/1 318,000 Acquired treasury stock for cash, 8/1 162,000 2-for-1 stock split, 10/1 Compute earnings per share. (Round answer to 2 decimal places, e.g. $2.55.) Earnings per share $enter earnings per share rounded to 2 decimal places
Answer #1

Weighted average share outstanding

Date
1/1 (816000*2*4/12) = 544000
5/1 1134000*2*3/12 = 567000
8/1 972000*2*2/12 = 324000
10/1 1944000*3/12 = 486000
Total 1921000

EPS = (2550000-408000)/1921000 = 1.12 per share